Okay...
I have 4 questions.
1. 6q=6(4q+1) What is "q" equal to?
2. 7(t-2)-t=4 What is "t" equal to?
3. 6w+4=4w+1 What is "w" equal to?
4. 13s=2s-66 What is "s" equal to?
I really don't get this and it is really hard. I need an explanation for all please and thank you very much.

Answer by ewatrrr(24785)


Hi,
Note: These questions all can be answered by following the steps needed for each
For Example:
1) removing the parenthesis
2) combining like terms
3) isolating the variable on one side of the equation
always keep in mind: Any operatioms done, need to done on both sides of the equation
.
6q=6(4q+1)
removing the parenthesis
6q= 24q+6
isolating the variable on one side of the equation
-6 = 18q
-1/3 = q
.
7(t-2)-t=4
removing the parenthesis
7t-14-t=4
combining like terms
6t = 28
t = 4.667
.
6w+4=4w+1
combining like terms
2w = -3
w = -3/2
.
13s=2s-66
combining like terms
11s = -66
s = -6